# Disordered Dynamics in High Dimensions: Connections to Random Matrices and Machine Learning

### v1

@@ 1 Introduction
In many areas of physics and applied mathematics, one encounters high dimensional dynamical systems which depend on some source of randomness. In statistics or machine learning theory, the randomness could come from randomly sampled data points krogh1990dynamics; advani2020high; montanari2025dynamical; agoritsas2018out, stochastic gradient noise mignacco2020dynamical; mignacco2022effective, or from the initialization of model parameters mei2022generalization; adlam2020neural. In physics, randomness in the interactions of a high dimensional system, such as a spin glass, can generate rich high dimensional dynamics de1978dynamics; sompolinsky1981dynamic; berthier2011theoretical. This idea has also been pursued in theoretical neuroscience, where randomly connected recurrent neural networks have been analyzed for decades as simple solvable models of high dimensional chaotic dynamics sompolinsky1988chaos; vogels2005neural; helias2020statistical. Many works in theoretical ecology have analyzed the population dynamics of interacting species with complex cross-species interactions may1972will; cui2021diverse; blumenthal2024phase; cui2024houches.
@@
  • Gradient Descent/Flow Dynamics of Linear Regression. We will consider linear regression on random covariates with $P$ data points in $D$ dimensions with $P=\alpha D$ krogh1990dynamics. For isotropic covariates, the DMFT equations will encode the Marchenko-Pastur Law advani2020high.
  • Gradient Descent for Kernel Regression. We next examine gradient flow on structured regression problems such as kernel methods. As an example, we will describe “stage like” multiple descent structures in problems with large numbers of degenerate eigenvalues lu2025equivalence; xiao2022precise; hu2024asymptotics as well as powerlaw behavior for powerlaw eigenspectra bordelon2020spectrum; paquette20244+; bordelon2024dynamical.
  • Random Feature Models. Instead of one random matrix from the randomly sampled training data, we then introduce another random matrix representing a feature projection mei2022generalization. In this case, we show that the spectrum of the matrix which governs the linear dynamics of the test loss is insufficient to characterize the loss dynamics. Rather the DMFT two-point correlation function can capture interesting non-monotonic behavior in the test loss near the interpolation threshold where parameters and data are equal.
@@
  • Non-Hermitian Examples. We provide a simple DMFT recipe to compute spectra for non-Hermitian matrices with complex spectra by studying long-time behavior of a Hermitianized flow. We use this to derive the circular law girko1985circular; feinberg9703118non; cui2024elementary & and the spectrum of a diagonally modulated asymmetric random matrix, which is related to the Jacobian of random recurrent neural networks krishnamurthy2022theory.
  • Beyond Linear Dynamics. We provide simple examples of systems which are not described by a linear dynamics where a (Gaussian) DMFT still provides the exact asymptotic description of the system. We will first illustrate a toy example of (Anti-)Hebbian dynamics on an initially GOE recurrent network before using this theory to describe multilayer linear networks in a proportional scaling limit where data, input dimension, and width are all comparable bordelon_deep_linear.
@@ 2 Disordered Linear Dynamical Systems
The trace of the resolvent, $\mathcal{R}(\omega)=\frac{1}{N}\text{Tr}\left[i\omega+\bm{M}\right]^{-1}$ is often termed the Stieltjes transform, is a central object in the study of random matrices potters2020first. For matrices $\bm{M}$ with real spectra, the eigenvalue density $\rho(\lambda)$ can be obtained from the Sokhotski–Plemelj formula
$\rho(\lambda)=\lim_{\epsilon\to 0}\frac{1}{\pi}\Im\ \mathcal{R}(\omega)|_{\omega=i\lambda-\epsilon}.$ (5)
@@ 3.2 Path Integral Approach
where the average is taken over the distribution of the random variables $\bm{h}(t)$ induced by the random couplings $\bm{M}$. While moments of the distribution for $\bm{h}$ can be obtained through derivatives of $Z$ with respect to $\bm{\zeta}$, it often suffices to analyze $Z$ formally for $\bm{\zeta}=0$, where $Z=1$ de1978dynamics; helias2020statistical; crisanti2018path, which we explain in Appendix A. Starting from an integral representation of the Dirac delta function $\delta(z)=\int\frac{d\hat{z}}{2\pi}\ \exp(i\hat{z}z)$, we can write a formal path integral that enforces the dynamics for $\bm{h}(t)$, yielding
@@ 3.3 Anti-Symmetrized Version
The oscillatory version of this system, where $\frac{d}{dt}\bm{h}(t)=\bm{M}\bm{h}(t)$ where $\bm{M}=\frac{1}{\sqrt{2N}}(\bm{A}-\bm{A}^{\top})$ where $A_{ij}\sim\mathcal{N}(0,1)$ was analyzed with DMFT in another recent note blumenthal2025building. In this case, one arrives at very similar DMFT equations modified by flipping the sign of the term involving the response function
@@ 4.1 Bipartite Cavity Analysis
For the regression problem, we introduced two sets of variables $\{\bm{h}(t),\bm{\Delta}(t)\}$ which are both linearly related through the matrix $\bm{\Psi}$. We can therefore perform two-steps of a cavity argument, one for each of these variables (see clark2025simplified for application of this idea for several static settings). This proceeds in two steps which are visualized in Figure 3. First, we consider adding a (N+1)(N+1)st feature dimension which leads to a new variable $h_{0}(t)$, which leads to new couplings to the original $P$ traning errors (red lines). Next, we consider adding the (P+1)(P+1)st data point, which couples to all $N$ original features (purple lines).
@@ 4.2 Path Integral Approach
$h(\omega)=\mathcal{R}_{h}(\omega)\left[w^{*}+u_{h}(\omega)\right]\ ,\ i\omega\ \mathcal{R}_{h}(\omega)=-1+\frac{\mathcal{R}_{h}(\omega)}{1-\alpha^{-1}\mathcal{R}_{h}(\omega)}$ (43)
@@
$\mathcal{R}_{h}(\omega)=\frac{\alpha}{2(i\omega)}\left[(1+i\omega-\alpha^{-1})+\sqrt{(1+i\omega-\alpha^{-1})^{2}+4(i\omega)\alpha^{-1}}\right]$ (44)
Again evaluating at $\omega=i\lambda-\epsilon$, taking the $\epsilon\to 0$ limit and using the fact that, the eigenvalue density $\rho(\lambda)$ is recovered
$\rho(\lambda)=\frac{1}{\pi}\Im\mathcal{R}(i\lambda-\epsilon)=\frac{1}{2\pi\lambda}\sqrt{\left[4\lambda\alpha^{-1}-(\alpha^{-1}-1+\lambda)^{2}\right]_{+}}+[1-\alpha]_{+}\delta(\lambda)$ (45)
@@ 5 Structured Covariates and Kernel Methods
The large time limit of these DMFT equations recover static computations for the test loss of the final predictor bordelon2020spectrum; canatar2021spectral; loureiro2021learning.
@@ 5.1 Gapped and Degenerate Spectra: Multiple Stages in Time and Data
In this section, we explore the case of a spectrum that consists of stages of a large number of degenerate eigenvalues. One commonly studied setting which generates this is dot-product kernels for spherically symmetric data in $D\gg 1$ dimensions mei2022generalization; bordelon2020spectrum; canatar2021spectral. In this case, eigenvalues $\lambda_{k}$ are all equal for the orthogonal degree $k$ polynomials (Hermite polynomials or spherical harmonics) which carry multiplicity $\mathcal{N}_{k}$ which scales polynomially with the dimension $D$ in the following way (for $D\gg k$)
@@ 5.2 Power Laws
where the exponents $\chi\nu$ and $\nu$ are known as the source and capacity exponents respectively cui2021generalization; bordelon2024dynamical. The solution to $\mathcal{R}_{\Delta}(\omega)$ can be approximated by an early time (high frequency) and late time (low frequency) expansion in $P(i\omega)^{\nu}$
@@
These equations indicate that the early stage of the dynamics, the loss dynamics along the $k$-th population eigenmode obey $\mathcal{H}_{k}(\tau)\approx e^{-\lambda_{k}\tau}+\mathcal{O}(P^{-1})$ for small times $\tau$ and converge to $\lim_{\tau\to\infty}\mathcal{H}_{k}(\tau)\approx\frac{1}{1+(P/k)^{\nu}}$ at late time. This indicates the intuitive fact that if $P\gg k$ this eigenmode will be learned effectively after sufficient training, while for $P\ll k$ this eigenmode cannot be resolved at this sample size $P$. As a consequence the loss can be roughly approximated as a combination of powerlaws (motivated by hoffmann2022training)
@@
in the sense that this expression captures the $\lim_{t\to\infty}\mathcal{L}\sim P^{-\nu\chi}$ and $\lim_{N\to\infty}\mathcal{L}\sim t^{-\chi}$. This is the same functional form explored in “Chinchilla” neural scaling laws hoffmann2022training; paquette20244+. Additional terms can be incorporated which capture the dynamical effect of variance from sampling the random dataset (which are mixed terms involving both finite $t,P$).
@@ 6.1 Online Stochastic Gradient Descent on Structured Random Features
We plot this discrete time solution against SGD simulations in Figure 9. Additional simulations of this model can be found in bordelon2024dynamical. A continuous time approximation of the above dynamics paquette2021sgd; paquette20244+; atanasov2025two; mignacco2022effective; mignacco2020dynamical.
@@ 7 Symmetric and Asymmetric Free Products
In this section, we move beyond Gaussian random matrices and consider general free products potters2020first; atanasov2024scaling; atanasov2025two; bordelon2025theory. Consider two $N\times N$ matrices $\bm{A}$ and $\bm{B}$ and define the two types of free products
@@ 7.2.1 Asymmetric Case
$\bm{h}(\omega)\bm{h}(\omega^{\prime})^{\top}\simeq\left(i\omega+\Psi(\omega)\bm{A}\right)^{-1}\left[\bm{h}_{0}\bm{h}_{0}^{\top}-\Psi_{\chi\chi}(\omega,\omega^{\prime})\bm{I}\right]\left(i\omega+\Psi(\omega)\bm{A}\right)^{-1}$ (121)
@@ 7.2.2 Symmetrized Case
$\bm{h}(\omega)\bm{h}(\omega^{\prime})^{\top}\simeq\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}\left[\bm{h}_{0}\bm{h}_{0}^{\top}-\Psi_{\chi\chi}(\omega,\omega^{\prime})\bm{A}\right]\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1},$ (124)
@@
$\Sigma_{hh}(\omega,\omega^{\prime})$ $=\frac{1}{1-\Gamma(\omega,\omega^{\prime})}\text{tr}\bm{A}\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}\bm{h}_{0}\bm{h}_{0}^{\top}\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}$ (126)
@@ 8 Non-Hermitian Systems with Complex Spectra
To counteract this problem, we can consider a dynamical system driven by a hermitian matrix which relaxes to a fixed point that depends on the resolvent of the matrix girko1985circular; feinberg9703118non; baron2022eigenvalues; cui2024elementary. We consider the following flows
@@
from which the spectral density in the complex plane at point $z\in\mathbb{C}$ can be obtained [^7] We use the definition $\partial_{z^{\star}}=\frac{1}{2}(\partial_{x}+i\partial_{y})$ for $z=x+iy$ so that $\partial_{z^{\star}}z^{\star}=1$ and $\partial_{z^{\star}}\ \frac{1}{z}=\delta(z)$.. Fortunately, this Hermitianized dynamical system does generate response functions in its associated DMFT equations that enable computation of the complex spectrum from the long time limit  [^8] For readers familiar with Hermitianization, often a regulator is introduced when computing the resolvent of the Hermitianized matrix sommers1988spectrum. In our setup, the finite time plays the role of the regulator, with the long-time limit $t\to\infty$ ($i\omega\to 0$) giving the final result..
@@ 8.1 Ginibre Matrices and the Circular Law
$\frac{d}{dt}h_{0}(t)=-\xi_{2}(t)-\int dt^{\prime}\ R_{1}(t,t^{\prime})h_{0}(t^{\prime})+z^{\star}h_{1}(t)$ (146)
$h_{1}(t)=\xi_{1}(t)+\int dt^{\prime}\ \mathcal{R}_{0,2}(t,t^{\prime})h_{1}(t^{\prime})-zh_{0}(t)+b$ (147)
Defining $\mathcal{H}(t,t^{\prime})=-R_{0,2}(t,t^{\prime})$ and taking a Fourier transform, we find
@@
This recovers the circular law, where the eigenvalue density is uniform in a unit disk in the complex plane ginibre1965statistical; sommers1988spectrum.

### v2

@@ 1 Introduction
In many areas of physics and applied mathematics, one encounters high dimensional dynamical systems which depend on some source of randomness. In statistics or machine learning theory, the randomness could come from randomly sampled data points Krogh and Hertz (1990); Advani et al. (2020); Montanari and Urbani (2025); Agoritsas et al. (2018), stochastic gradient noise Mignacco et al. (2020); Mignacco and Urbani (2022), or from the initialization of model parameters Mei and Montanari (2022); Adlam and Pennington (2020). In physics, randomness in the interactions of a high dimensional system, such as a spin glass, can generate rich high dimensional dynamics De Dominicis (1978); Sompolinsky and Zippelius (1981); Berthier and Biroli (2011). This idea has also been pursued in theoretical neuroscience, where randomly connected recurrent neural networks have been analyzed for decades as simple solvable models of high dimensional chaotic dynamics Sompolinsky et al. (1988); Vogels et al. (2005); Helias and Dahmen (2020). Many works in theoretical ecology have analyzed the population dynamics of interacting species with complex cross-species interactions May (1972); Cui et al. (2021b); Blumenthal et al. (2024); Cui et al. (2024a).
@@
  • Gradient Descent/Flow Dynamics of Linear Regression. We will consider linear regression on random covariates with $P$ data points in $D$ dimensions with $P=\alpha D$ Krogh and Hertz (1990). For isotropic covariates, the DMFT equations will encode the Marchenko-Pastur Law Advani et al. (2020).
  • Gradient Descent for Kernel Regression. We next examine gradient flow on structured regression problems such as kernel methods. As an example, we will describe “stage like” multiple descent structures in problems with large numbers of degenerate eigenvalues Lu and Yau (2025); Xiao et al. (2022); Hu et al. (2024) as well as powerlaw behavior for powerlaw eigenspectra Bordelon et al. (2020); Paquette et al. (2024); Bordelon et al. (2024).
  • Random Feature Models. Instead of one random matrix from the randomly sampled training data, we then introduce another random matrix representing a feature projection Mei and Montanari (2022). In this case, we show that the spectrum of the matrix which governs the linear dynamics of the test loss is insufficient to characterize the loss dynamics. Rather the DMFT two-point correlation function can capture interesting non-monotonic behavior in the test loss near the interpolation threshold where parameters and data are equal.
@@
  • Non-Hermitian Examples. We provide a simple DMFT recipe to compute spectra for non-Hermitian matrices with complex spectra by studying long-time behavior of a Hermitianized flow. We use this to derive the circular law Girko (1985); Feinberg and Zee ; Cui et al. (2024b) & and the spectrum of a diagonally modulated asymmetric random matrix, which is related to the Jacobian of random recurrent neural networks Krishnamurthy et al. (2022).
  • Beyond Linear Dynamics. We provide simple examples of systems which are not described by a linear dynamics where a (Gaussian) DMFT still provides the exact asymptotic description of the system. We will first illustrate a toy example of (Anti-)Hebbian dynamics on an initially GOE recurrent network before using this theory to describe multilayer linear networks in a proportional scaling limit where data, input dimension, and width are all comparable Bordelon and Pehlevan (2025).
@@ 2 Disordered Linear Dynamical Systems
The trace of the resolvent, $\mathcal{R}(\omega)=\frac{1}{N}\text{Tr}\left[i\omega+\bm{M}\right]^{-1}$ is often termed the Stieltjes transform, is a central object in the study of random matrices Potters and Bouchaud (2020). For matrices $\bm{M}$ with real spectra, the eigenvalue density $\rho(\lambda)$ can be obtained from the Sokhotski–Plemelj formula
$\rho(\lambda)=\lim_{\epsilon\to 0}\frac{1}{\pi}\Im\ \mathcal{R}(\omega)|_{i\omega=-\lambda-i\epsilon}.$ (5)
@@ 3.2 Path Integral Approach
where the average is taken over the distribution of the random variables $\bm{h}(t)$ induced by the random couplings $\bm{M}$. While moments of the distribution for $\bm{h}$ can be obtained through derivatives of $Z$ with respect to $\bm{\zeta}$, it often suffices to analyze $Z$ formally for $\bm{\zeta}=0$, where $Z=1$ De Dominicis (1978); Helias and Dahmen (2020); Crisanti and Sompolinsky (2018), which we explain in Appendix A. Starting from an integral representation of the Dirac delta function $\delta(z)=\int\frac{d\hat{z}}{2\pi}\ \exp(i\hat{z}z)$, we can write a formal path integral that enforces the dynamics for $\bm{h}(t)$, yielding
@@ 3.3 Anti-Symmetrized Version
The oscillatory version of this system, where $\frac{d}{dt}\bm{h}(t)=\bm{M}\bm{h}(t)$ where $\bm{M}=\frac{1}{\sqrt{2N}}(\bm{A}-\bm{A}^{\top})$ where $A_{ij}\sim\mathcal{N}(0,1)$ was analyzed with DMFT in another recent note Blumenthal (2025). In this case, one arrives at very similar DMFT equations modified by flipping the sign of the term involving the response function
@@ 4.1 Bipartite Cavity Analysis
For the regression problem, we introduced two sets of variables $\{\bm{h}(t),\bm{\Delta}(t)\}$ which are both linearly related through the matrix $\bm{\Psi}$. We can therefore perform two-steps of a cavity argument, one for each of these variables (see Clark and Sompolinsky (2025) for application of this idea for several static settings). This proceeds in two steps which are visualized in Figure 3. First, we consider adding a (N+1)(N+1)st feature dimension which leads to a new variable $h_{0}(t)$, which leads to new couplings to the original $P$ traning errors (red lines). Next, we consider adding the (P+1)(P+1)st data point, which couples to all $N$ original features (purple lines).
@@ 4.2 Path Integral Approach
$h(\omega)=\mathcal{R}_{h}(\omega)\left[\beta_{\star}+u_{h}(\omega)\right]\ ,\ i\omega\ \mathcal{R}_{h}(\omega)=1-\frac{\mathcal{R}_{h}(\omega)}{1+\alpha^{-1}\mathcal{R}_{h}(\omega)}$ (43)
@@
$\mathcal{R}_{h}(\omega)=\frac{\alpha}{2(i\omega)}\left[-(1+i\omega-\alpha^{-1})+\sqrt{(1+i\omega-\alpha^{-1})^{2}+4(i\omega)\alpha^{-1}}\right]$ (44)
Again evaluating at $i\omega=-\lambda-i\epsilon$, taking the $\epsilon\to 0$ limit and using the fact that, the eigenvalue density $\rho(\lambda)$ is recovered
$\rho(\lambda)=\frac{1}{\pi}\Im\ \mathcal{R}(\omega)|_{i\omega=-\lambda-i\epsilon}=\frac{1}{2\pi\lambda}\sqrt{\left[4\lambda\alpha^{-1}-(\alpha^{-1}-1+\lambda)^{2}\right]_{+}}+[1-\alpha]_{+}\delta(\lambda)$ (45)
@@ 5 Structured Covariates and Kernel Methods
The large time limit of these DMFT equations recover static computations for the test loss of the final predictor Bordelon et al. (2020); Canatar et al. (2021); Loureiro et al. (2021).
@@ 5.1 Gapped and Degenerate Spectra: Multiple Stages in Time and Data
In this section, we explore the case of a spectrum that consists of stages of a large number of degenerate eigenvalues. One commonly studied setting which generates this is dot-product kernels for spherically symmetric data in $D\gg 1$ dimensions Mei and Montanari (2022); Bordelon et al. (2020); Canatar et al. (2021). In this case, eigenvalues $\lambda_{k}$ are all equal for the orthogonal degree $k$ polynomials (Hermite polynomials or spherical harmonics) which carry multiplicity $\mathcal{N}_{k}$ which scales polynomially with the dimension $D$ in the following way (for $D\gg k$)
@@ 5.2 Power Laws
where the exponents $\chi\nu$ and $\nu$ are known as the source and capacity exponents respectively Cui et al. (2021a); Bordelon et al. (2024). The solution to $\mathcal{R}_{\Delta}(\omega)$ can be approximated by an early time (high frequency) and late time (low frequency) expansion in $P(i\omega)^{\nu}$
@@
These equations indicate that the early stage of the dynamics, the loss dynamics along the $k$-th population eigenmode obey $\mathcal{H}_{k}(\tau)\approx e^{-\lambda_{k}\tau}+\mathcal{O}(P^{-1})$ for small times $\tau$ and converge to $\lim_{\tau\to\infty}\mathcal{H}_{k}(\tau)\approx\frac{1}{1+(P/k)^{\nu}}$ at late time. This indicates the intuitive fact that if $P\gg k$ this eigenmode will be learned effectively after sufficient training, while for $P\ll k$ this eigenmode cannot be resolved at this sample size $P$. As a consequence the loss can be roughly approximated as a combination of powerlaws (motivated by Hoffmann et al. (2022))
@@
in the sense that this expression captures the $\lim_{t\to\infty}\mathcal{L}\sim P^{-\nu\chi}$ and $\lim_{N\to\infty}\mathcal{L}\sim t^{-\chi}$. This is the same functional form explored in “Chinchilla” neural scaling laws Hoffmann et al. (2022); Paquette et al. (2024). Additional terms can be incorporated which capture the dynamical effect of variance from sampling the random dataset (which are mixed terms involving both finite $t,P$).
@@ 6.1 Online Stochastic Gradient Descent on Structured Random Features
We plot this discrete time solution against SGD simulations in Figure 9. Additional simulations of this model can be found in Bordelon et al. (2024). A continuous time approximation of the above dynamics Paquette et al. (2021, 2024); Atanasov et al. (2025); Mignacco and Urbani (2022); Mignacco et al. (2020).
@@ 7 Symmetric and Asymmetric Free Products
In this section, we move beyond Gaussian random matrices and consider general free products Potters and Bouchaud (2020); Atanasov et al. (2024, 2025); Bordelon et al. (2025). Consider two $N\times N$ matrices $\bm{A}$ and $\bm{B}$ and define the two types of free products
@@ 7.2.1 Asymmetric Case
$\bm{h}(\omega)\bm{h}(\omega^{\prime})^{\top}\simeq\left(i\omega+\Psi(\omega)\bm{A}\right)^{-1}\left[\bm{h}_{0}\bm{h}_{0}^{\top}-\Psi_{\chi\chi}(\omega,\omega^{\prime})\bm{I}\right]\left(i\omega^{\prime}+\Psi(\omega^{\prime})\bm{A}\right)^{-1}$ (121)
@@ 7.2.2 Symmetrized Case
$\bm{h}(\omega)\bm{h}(\omega^{\prime})^{\top}\simeq\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}\left[\bm{h}_{0}\bm{h}_{0}^{\top}-\Psi_{\chi\chi}(\omega,\omega^{\prime})\bm{A}\right]\left(i\omega^{\prime}+\Psi_{h\chi}(\omega^{\prime})\bm{A}\right)^{-1},$ (124)
@@
$\Sigma_{hh}(\omega,\omega^{\prime})$ $=\frac{1}{1-\Gamma(\omega,\omega^{\prime})}\text{tr}\bm{A}\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}\bm{h}_{0}\bm{h}_{0}^{\top}\left(i\omega^{\prime}+\Psi_{h\chi}(\omega^{\prime})\bm{A}\right)^{-1}$ (126)
@@ 8 Non-Hermitian Systems with Complex Spectra
To counteract this problem, we can consider a dynamical system driven by a hermitian matrix which relaxes to a fixed point that depends on the resolvent of the matrix Girko (1985); Feinberg and Zee ; Baron et al. (2022); Cui et al. (2024b). We consider the following flows
@@
from which the spectral density in the complex plane at point $z\in\mathbb{C}$ can be obtained [^7] We use the definition $\partial_{z^{\star}}=\frac{1}{2}(\partial_{x}+i\partial_{y})$ for $z=x+iy$ so that $\partial_{z^{\star}}z^{\star}=1$ and $\partial_{z^{\star}}\ \frac{1}{z}=\delta(z)$.. Fortunately, this Hermitianized dynamical system does generate response functions in its associated DMFT equations that enable computation of the complex spectrum from the long time limit  [^8] For readers familiar with Hermitianization, often a regulator is introduced when computing the resolvent of the Hermitianized matrix Sommers et al. (1988). In our setup, the finite time plays the role of the regulator, with the long-time limit $t\to\infty$ ($i\omega\to 0$) giving the final result..
@@ 8.1 Ginibre Matrices and the Circular Law
$\frac{d}{dt}h_{0}(t)=-h^{2}(t)\ ,\ h^{2}(t)=\xi_{2}(t)+\int dt^{\prime}\ R_{1}(t,t^{\prime})h_{0}(t^{\prime})-z^{\star}h_{1}(t)$ (146)
$h_{1}(t)=\xi_{1}(t)+\int dt^{\prime}\ R_{0,2}(t,t^{\prime})h_{1}(t^{\prime})-zh_{0}(t)+b$ (147)
where $R_{1}(t,t^{\prime})=\frac{\partial h_{1}(t)}{\partial\xi_{1}(t^{\prime})}$ and $R_{0,2}(t,t^{\prime})=\frac{\partial h_{0}(t)}{\partial\xi_{2}(t^{\prime})}$ are the response functions that arise from the DMFT equations. Defining $\mathcal{H}(t,t^{\prime})=-R_{0,2}(t,t^{\prime})$ and taking a Fourier transform, we find
@@
This recovers the circular law, where the eigenvalue density is uniform in a unit disk in the complex plane Ginibre (1965); Sommers et al. (1988).